40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 12 20:15:22 2016, Lydia, Update, SUS, Diagonalization in air ITMY_fit.jpgBS_diag.jpg
    Reply  Tue Sep 13 19:18:43 2016, Lydia, Update, SUS, Diagonalization in air ITMY_diag.jpg38.png
       Reply  Wed Sep 14 20:27:03 2016, Lydia, Update, SUS, Diagonalization ETMY_diag.png
          Reply  Thu Sep 15 18:37:20 2016, Lydia, Update, SUS, Diagonalization ETMY_osemspec.pngITMY_osemspec.png
             Reply  Fri Sep 16 19:14:27 2016, Lydia, Update, SUS, Diagonalization ETMY_yawdrivedefault.pngETMY_yawdrivenew.pngETMY_yawdriveold.png57.png
Message ID: 12490     Entry time: Tue Sep 13 19:18:43 2016     In reply to: 12484     Reply to this: 12495
Author: Lydia 
Type: Update 
Category: SUS 
Subject: Diagonalization in air 

[Lydia, Teng]

We continued to work on the diagonalization scripts today and devised a way of choosing starting parameters that seems to work much better, and is easier to use, than tuning up to 15 parameters by hand per optic.

  • As before, the spectrum for each dof is estimated by using the "ideal" input matrix.
  • The starting guess for the peak frequency for each dof is the bin which achieves the maximum value of the spectrum between 0.4 and 1.5 Hz.
  • If another dof has a higher value at that frequency, the next highest peak is used. (Sometimes, for example, the peak in PIT at the POS frequency is stronger than the real POS peak!)
  • The peak height is initially guessed to be the spectrum value at the initial frequncy guess.
  • The width paramter Q can still be read from a file, but for all the times we tried, the peaks were found successfully if Q was initially guessed to be 300, so there might be no need to do this.
  • Spectra should still be examined to make sure the results make sense, and once we look at free swinging data in vacuum, we should compare the frequency results to the wiki values.
  • Reasonably good matrix values are saved to peakFit/inMats/1157630417. We got good diagonalization results for all but ITMY (see below). The values used for damping have not been overwritten.

We still noticed phase problems with ITMY, which appear to be preventing good diagonalization (See Attachment 1). Almost every degree of freedom has a significant imaginary part in the sensing matrix. We looked at the phases of the cross spectra in DDT and saw that indeed, the OSEM signals do not have the appropriate relative phases at the peak frequencies, especially in PIT and YAW (see Attachment 2: the phase at the peak is about 30 degrees when it should be 180). These phases are different for data takes ~24 hours apart, but are still wrong. We also looked at this information for ETMY and saw the correct behavior. We temporarily moved the pitch and yaw sliders for ITMY and looked at the OSEM response on a striptool, and the signals moved in the expected way. Can anyone suggest a reason why this would be happening? Is there another stretch of data (besides this past weekend) which would be good to compare to?

 

Attachment 1: ITMY_diag.jpg  202 kB  | Hide | Hide all
ITMY_diag.jpg
Attachment 2: 38.png  28 kB  | Hide | Hide all
38.png
ELOG V3.1.3-